OpenStudy (anonymous):

The population of Guatemala in 2000 was 12.7 million. 1. Assuming exponential growth, what would be the size of the population after time t (measured in years after 2000) if the population was 30 million in 2020? Answer (in millions): P(t)= 2. Assuming exponential growth, what would be the size of the population after time t (measured in years after 2000) if the population was 30 million in 2075? Answer (in millions): P(t)=

OpenStudy (anonymous):

@campbell_st @phi

OpenStudy (anonymous):

@Hero @bahrom7893 @precal

OpenStudy (anonymous):

whats the difference between the old amount and the initial amount

hero (hero):

\[A = A_0e^{kt}\] \[A = \text{new amount} \\A_0 = \text{initial amount} \\k=\text{percent growth or decay rate} \\t=\text{time(usually in years)}\]

OpenStudy (anonymous):

so how would i find p(t)

hero (hero):

A = P(t)

hero (hero):

\[P(t) = P_ie^{kt}\]

OpenStudy (anonymous):

what would k be

hero (hero):

You have to find k

OpenStudy (anonymous):

how

OpenStudy (anonymous):

@zepdrix @robtobey @radar @tcarroll010

hero (hero):

@muzzammil.raza

OpenStudy (anonymous):

yeah?

hero (hero):

Just plug in the numbers and solve

OpenStudy (anonymous):

12700000e^(20k)?

hero (hero):

I wonder if they want P(t) or P(20)

hero (hero):

P(t) is just a general formula. P(20) is already given.

OpenStudy (anonymous):

they want p(t)

hero (hero):

If they want P(t) as it relates to the first question, then you have to find the values of \(P_0\) and \(k\)

hero (hero):

Plug those values into the formula \[P(t) = P_0e^{kt}\]

hero (hero):

You find k by inputting all the values, then solving for k

OpenStudy (anonymous):

inputting all the values into what

hero (hero):

Into the formula bro.

hero (hero):

Okay, I'll show you how to really do it.

hero (hero):

They give you all the information you need to find k

hero (hero):

They already given you \(P_0\)

hero (hero):

\(P_0 = 12.7\)

hero (hero):

\(P(20) = 30\)

hero (hero):

and t = 20

OpenStudy (anonymous):

ok so k=30 for the first one

hero (hero):

?

OpenStudy (anonymous):

i'm lost

hero (hero):

Bro, are you making stuff up

OpenStudy (anonymous):

no i am trying to find out what u r doing

hero (hero):

\(30 = 12.7e^{k(20)}\)

hero (hero):

I never mentioned anything about the value of k

hero (hero):

Because it is what you have to find

hero (hero):

\[\frac{30}{12.7} = e^{20k}\]

OpenStudy (anonymous):

12.7 million right

hero (hero):

Now take the natural log of both sides: \[\ln\left(\frac{30}{12.7}\right) = \ln{e^{20k}}\]

hero (hero):

You leave off the million. It would be awkward to include it. We already know we're dealing with millions of people

hero (hero):

Continue solving for k.

OpenStudy (anonymous):

i am getting a negative number for ln(30/12.7)

hero (hero):

It's not negative.

hero (hero):

Try again

OpenStudy (anonymous):

now i am getting .85959595959595...

hero (hero):

Good. Now what?

OpenStudy (anonymous):

.04298...

hero (hero):

That's the value of k

OpenStudy (anonymous):

so whats the answer for the question

hero (hero):

Now rewrite the general equation of P(t) but only inlude the values for \(P_0\) and k

hero (hero):

\[P(t) = 12.7e^{0.043t}\]

OpenStudy (anonymous):

so thats the answer

hero (hero):

That's your P(t) for the first one. You're on your own for the second one.

OpenStudy (anonymous):

please help me with this one i don't know how to do it

hero (hero):

You know how to do it. I showed you how to do the first one. The steps for doing the second one is the same as the first.

OpenStudy (anonymous):

i still don't know how to do it. i am basically mentaly challenged when it comes to math

hero (hero):

Everyone is mentally challenged when it comes to math. But if you keep telling yourself you can't do something, then you won't be able to do it. But imagine what will happen if you keep telling yourself that you CAN do it.

OpenStudy (anonymous):

dr.phil please just help me

OpenStudy (anonymous):

i alredy tried to do it twice and got it wrong

hero (hero):

I'm sorry to hear that.